OpenStudy (anonymous):

The second angle of a triangle garden is four times as large as the first. The third and is 45 degrees less then the sum of the other two angles. Find the measures of each angle.

OpenStudy (anonymous):

is it x=40.5

OpenStudy (anonymous):

let x be the angle of the first and y be the the second and z be the third y=4x z=x+y-45 =x+4x-45 =5x-45 Now x+y+z=180 x+4x+5x-45=180 10x=225 x=22.5 y=90 z=67.5

OpenStudy (mertsj):



OpenStudy (mertsj):

Just like Zed said.